Two angles are supplementary. One angle has a measure that is five less than four times the other . What is the measure of the larger angle?
Answer by nerdybill(7384)

Two angles are supplementary. One angle has a measure that is five less than four times the other . What is the measure of the larger angle?
.
When two angles are supplementary, the sum of the two angles is 180 degrees.
.
Let x = one angle
then from: "One angle has a measure that is five less than four times the other"
4x-5 = second angle
.
x + 4x-5 = 180
5x-5 = 180
x-1 = 36
x = 37 deg
.
second angle:
4x-5 = 4(37)-5 = 148-5= 143 degrees (larger angle)